40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Aug 18 23:59:53 2017, gautam, Update, ALS, X Arm ALS lock 
    Reply  Sat Aug 19 01:35:08 2017, ericq, Update, ALS, X Arm ALS lock ALS_comparison.pdf
    Reply  Mon Aug 21 23:38:55 2017, gautam, Update, ALS, ALS out-of-loop noise ALS_21082017.pdf
       Reply  Tue Aug 22 02:19:11 2017, gautam, Update, ALS, ALS OLTFs ALS_OLTFs.pdfALS_OLTF_Aug2017.zip
          Reply  Tue Aug 22 23:27:14 2017, rana, Update, ALS, ALS OLTFs 
Message ID: 13238     Entry time: Tue Aug 22 02:19:11 2017     In reply to: 13237     Reply to this: 13244
Author: gautam 
Type: Update 
Category: ALS 
Subject: ALS OLTFs 

Attachment #1 shows the results of my measurements tonight (SR785 data in Attachment #2). Both loops have a UGF of ~10kHz, with ~55 degrees of phase margin.

Excitation was injected via SR560 at the PDH error point, amplitude was 35mV. According to the LED indicators on these boxes, the low frequency boost stages were ON. Gain knob of the X end PDH box was at 6.5, that of the Y end PDH box was at 4.9. I need to check the schematics to interpret these numbers. GV Edit: According to this elog, these numbers mean that the overall gain of the X end PDH box is approx. 25dB, while that of the Y end PDH box is approx. 15dB. I believe the Y end Lightwave NPRO has an actuator discriminant ~5MHz/V, while the X end Innolight is more like 1MHz/V.

Not sure what to make of the X PDH loop measurement being so much noisier than the Y end, I need to think about this.

More detailed analysis to follow.

Quote:

 

I am now going to measure the OLTFs of both green PDH loops to check that the overall loop gain is okay, and also check the measurement against EricQ's LISO model of the (modified) AUX green PDH servos. Results to follow.

 

Attachment 1: ALS_OLTFs.pdf  168 kB  | Hide | Hide all
ALS_OLTFs.pdf
Attachment 2: ALS_OLTF_Aug2017.zip  27 kB  Uploaded Tue Aug 22 03:19:50 2017
ELOG V3.1.3-